OpenStudy (anonymous):

If f(x)=(x-2)^2 and g(x)=x^2-2, what is the value of g(f[-2])?

OpenStudy (anonymous):

ooh this is easy

OpenStudy (anonymous):

please help? :)

OpenStudy (anonymous):

first solve for f(-2)= (-2-2)^2=-4^2 =16 so g(16)=16^2-2

OpenStudy (anonymous):

254

OpenStudy (anonymous):

oh wow. thanks though :p

OpenStudy (anonymous):

De nada